Unknown nuclear reaction proceeds by the hypothetical X --> Y + Z. At some beginning time, the activity of a sample is 1.9x1015 per second and produces 298Watts of power. After one year, the activity of the sample is 1.2x1015 per second. 1) If it is known that the mass of (Y+Z) is 123.456789012u, what is the mass of X (in u)? 2) What is the expected power output after 4 years from the beginning time?
Answer #1

1)
amount of energy released per per one reaction,

E = 298/(1.9*10^15)

= 1.5684210526*10^-13 J

The amount of mass converted to energy, m = E/c^2

= 1.5684210526*10^-13/(3*10^8)^2

= 1.7426900584*10^-30 kg

= 1.7426900584*10^-30/(1.66054*10^-27)

= 0.00104947189 u

so mass of X = mass of (X + Y) + m

= 123.456789012 + 0.00104947189

= 123.457838483 u <<<<<<<<<<-----------------------Answer

2) Expected power after 4 year, P_new = P_old*(1.2*10^15/(1.9*10^15))

= 298*1.2/1.9

= 188 W <<<<<<<<<<-----------------------Answer